Kolby’s Korndogs is looking at a new sausage system with an installed cost of $715,000. The asset qualifies for 100 percent bonus depreciation and can be scrapped for $97,000 at the end of the project’s 5-year life. The sausage system will save the firm $207,000 per year in pretax operating costs, and the system requires an initial investment in net working capital of $59,000. If the tax rate is 22 percent and the discount rate is 8 percent, what is the NPV of this project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Total initial investment = installed cost of new system + investment in net working capital = 715000 + 59000 = $774000

Since the asset qualities for 100% bonus depreciation, so

Depreciation in year 1 = 100% of Installed cost of new system = 100% x 715000 = $715000

Depreciation per year from year 2 to year 5 = 0

Incremental revenue = 0 , Incremental operating costs = -$207000

Operating cash flow for year 1 = (Incremental revenue - incremental operating costs - Depreciation)(1-tax rate) + Depreciation = [0 - (-207000) - 715000][1-22%] + 715000 = -508000 x 78% + 715000 = -396240 + 715000 = $318760

Operating cash flow per year from year 2 to year 5 = (Incremental revenue - incremental operating costs - Depreciation)(1-tax rate) + Depreciation = [0 - (-207000) - 0][1-22%] + 0 = 207000 x 78% + 0 = $161460

Book value of system at end of year 5 = Installed cost of system - Sum of depreciation for year 1 to 5 = 715000 - [715000 + 0 + 0 + 0 + 0] = 0

Terminal cash flow in year 5 = Salvage value at end of year 5 - Tax from gain on sale + Recovery of investment in net working capital = Salvage value at end of year 5 - Tax rate (Salvage value at end of year 5 - Book value at end of year 5) + Recovery of investment in net working capital = 97000 - 22%(97000 - 0) + 59000 = 97000 - 21340 + 59000 = 134660

NPV of the project = - Total initial investment + Sum of present value of operating cash flows for year 1 to year 5 + Present value of Terminal cash flow = -774000 + 318760/ (1+8%) + 161460 / (1+8%)2 + 161460 / (1+8%)3 + 161460 / (1+8%)4 + 161460 / (1+8%)5 + 134660 / (1+8%)5 = 295148.1481 + 138425.9259 + 128172.1536 + 118677.9200 + 109886.9629 + 91647.3333 = 107958.4431 = $107958.44 (rounded to two decimal places)

Hence NPV of the project = $107958.44